OpenStudy (anonymous):

squareroot of 2n squared - 121 = a pls help pls answer .

OpenStudy (anonymous):

What are you trying to find out?

OpenStudy (anonymous):

radical equation .

OpenStudy (anonymous):

So you have \[\sqrt{2n^2-121}=a\] Correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[n = \pm \sqrt{(a^2 + 121)/2}\]

OpenStudy (anonymous):

whats the solution and answer pls .

OpenStudy (anonymous):

what are you solving for? n? a? or are you factorizing?

OpenStudy (anonymous):

Can you write out the complete question?

OpenStudy (anonymous):

What kind of solution ?

OpenStudy (anonymous):

yes i am looking for n . squareroot of 2n^2 - n-=a

OpenStudy (anonymous):

a = -1/8 n = 1/4

OpenStudy (anonymous):

How did you get that answer Aron West?

OpenStudy (anonymous):

a = 2 m^2+m, n = -m, \[a = 2 m^2+m, n = -m, m \to \mathbb{Z}^+\]

OpenStudy (anonymous):

Sorry I don't follow, could you expand on this?

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

ALSO, \[n = 1/4(1 \pm \sqrt{8a +1})\]